An unbalanced coin shows heads 55% of the time.
(a) If the coin is tossed 12 times, what is the probability of heads on all tosses?

(b) Find the probability of at least one head in 12 tosses.

Answer by stanbon(75887)

An unbalanced coin shows heads 55% of the time.
(a) If the coin is tossed 12 times, what is the probability of heads on all tosses?
P(all heads) = 0.55^12 = 0.000766
---------------------
(b) Find the probability of at least one head in 12 tosses.
P(1<= x <=12) = 1 - P(x = 0) = 1 - 0.45^12 = 0.00003
-------------
Cheers,
Stan H.
-----------